A seamstress purchases 100 yds. of fabric. She can use it to make 20 skirts and 8 dresses. If she makes 14 dresses will she have enough fabric to make 10 skirts. How many yards of fabric are needed to make each?

Answer by stanbon(75887)

A seamstress purchases 100 yds. of fabric. She can use it to make 20 skirts and 8 dresses. If she makes 14 dresses will she have enough fabric to make 10 skirts. How many yards of fabric are needed to make each?
---------------------
Let "s" be # of yards needed to make a skirt.
Let "d" be # of yards needed to make a dress.
------------------------
EQUATIONS:
1st: 20s+ 8d=100
2nd: 10s+14d=100
------------------
Multiply 2nd by 2 to get:
3rd: 20s+28d=200
------------------
Subtract 1st from 3rd to solve for "d"
20d=100
d=5 (# of dresses to be made)
----------------
Substitute into 20s+8d=100 to solve for "s"
20s+8*5=100
20s = 60
s=3 (# of skirts be be made)
================
Cheers,
Stan H.